Danny's father is 4 times as old as Danny. Five years ago, Danny's father was 5 times older. How old is Danny?
I have guessed the answer correct but would like to be shown how to work it out, using the correct method. Thankyou.

Answer by lwsshak3(11628)

Danny's father is 4 times as old as Danny. Five years ago, Danny's father was 5 times older. How old is Danny?
I have guessed the answer correct but would like to be shown how to work it out, using the correct method.
***
at present:
let x=Danny's age
4x=Father's age
..
5 yrs ago
x-5=Danny's age
4x-5=Father's age
..
4x-5=5(x-5)
4x-5=5x-25
x=20
How old is Danny? 20